OpenStudy (anonymous):

Find the exact value of cos-1 square root of 3/2

OpenStudy (anonymous):

\[Let x=\cos^{-1} \frac{ \sqrt{3} }{ 2 },\cos x=\frac{ \sqrt{3} }{ 2 }=\cos \frac{ \pi }{ 6 },\]\[\cos \left( 2\pi-\frac{ \pi }{6 } \right)\] \[x=\frac{ \pi }{6 },\frac{ 11\pi }{6 } \inthe (0,2\pi)\]